# Eisenstein degeneration of Euler systems

### v1

@@ 4. Beilinson–Kato elements
The Kato Euler system (or Beilinson–Kato Euler system) is one of the most significant constructions in the study of $p$-adic methods in Number Theory. It has enabled the proof of new cases of the Bloch–Kato conjecture and provided a general approach to establish one divisibility in the Iwasawa main conjecture. More concretely, the Kato Euler system of [Kat04] can be attached to a general Coleman family ${\mathbf{f}}$, as discussed e.g.  by Hansen in [Han15] or by Benois and Büyükboduk in [BB22] and [BB24]; for the the case of Hida families, the theory is known since the work of Ochiai [Och03]. In this section, we begin by recalling the (two-variable) $p$-adic $L$-function of Kitagawa and Mazur, and later discuss the reciprocity law, emphasizing the role of the (one-variable) adjoint $p$-adic $L$-function.
@@ 4.1. The Kitagawa–Mazur $p$-adic $L$-function
With the previous notations, let $U$ be the open of the weight space over which the family is defined. Then we have an $\mathcal{O}(U)$-module $V({\mathbf{f}})^{*}$, free of rank 2.
We choose bases $\gamma^{\pm}_{{\mathbf{f}}}$ of the complex-conjugation eigenspaces $V({\mathbf{f}})^{(c=\pm 1)}$, which are free of rank one over $\mathcal{O}(U)$.
@@
$L_{p}({\mathbf{f}})\in\mathcal{O}(U)\mathop{\hat{\otimes}}\mathcal{H}_{\Gamma}$
@@
More precisely, the value at $(k,j)$, with $20\leqslant j\leqslant k-$ 2, interpolates the $L$-value $L(f_{k},1+j)$ up to appropriate factors, where $f$ is the weight $k$ specialization. See e.g. [FK24, §4] or [BB22] for the precise form of the interpolation property, which involves different kinds of factors.
@@ 4.2. The adjoint $p$-adic $L$-function
For this section, we mainly follow the works of Hida [Hid16] and Bellaïche [Bel21, §9], who constructed an adjoint $p$-adic $L$-function, using Kim’s scalar product on overconvergent modular symbols. See also [Mak23] for a detailed treatment of the topic.
@@
As pointed out in [Mak23], Bellaïche’s interpolation formula is only given at crystalline points of trivial nebentype, while in Thm. 2.3.7 of loc. cit. the author gives a much more general result.
@@ 4.3. Hida–Rankin vs Kitagawa–Mazur
This follows by a direct comparison of the critical values, as in [BD14, Thm. 3.4], noting that the Euler factors also match. Observe, however, that in this case we cannot discard the factors contributing to the adjoint, since they can be zero (and indeed, they vanish in the situations we want to study). ∎
@@ 4.4. Kato’s Euler system
Recall that we have chosen bases $\gamma^{\pm}_{{\mathbf{f}}}$ of the complex-conjugation eigenspaces $V({\mathbf{f}})^{(c=\pm 1)}$, which are free of rank one over $\mathcal{O}(U)$. For the moment, we assume that we are outside the Eisenstein case.
@@
$\mathcal{E}_{N}(x)=\prod_{\ell\mid N}(1-a_{\ell}({\mathbf{f}})\sigma_{\ell}^{-1})\in\Lambda_{U}.$
@@
More precisely, let $\mathcal{F}^{+}V({\mathbf{f}})$ denote the rank 1 unramified subrepresentation of $V({\mathbf{f}})$ over $\mathcal{O}(U)$ (which exists since ${\mathbf{f}}$ is ordinary); and let $\eta_{{\mathbf{f}}}\in\mathbf{D}_{\mathrm{cris}}(\mathcal{F}^{+}V({\mathbf{f}}))$ be the canonical vector constructed in [KLZ17], which is characterized by interpolating the classes $\eta_{f}$ of Section 2.3 for each classical specialization $f$ of ${\mathbf{f}}$. Then we have the following reciprocity law, where we have an extra contribution coming from the Euler factors at bad primes. For the following proposition, write $\mathcal{L}_{\mathcal{F}^{-}V({\mathbf{f}})^{*}}^{\operatorname{PR}}$ for the Perrin-Riou logarithm attached to $\mathcal{F}^{-}V({\mathbf{f}})^{*}$.
@@
The following explicit reciprocity law holds:
@@
This is Kato’s explicit reciprocity law in the framework of Coleman families, as developed e.g. in [BB22]. The proof when the Coleman family passes through a critical Eisenstein point works in the same way. ∎
This must be understood as an equality in two variables, with $L_{p}({\mathbf{f}},\chi)$ playing an auxiliary role. In many case, it is possible to invert the factor $L_{p}({\mathbf{f}},\chi)$ and normalize the Kato class dividing by it; however, in the Eisenstein case one must be more circumspect and we cannot do this normalization.
@@
In the setting where ${\mathbf{f}}$ is not a critical Coleman family, one usually renormalizes the classes to avoid dependence on the character $\chi$. More precisely, we may consider the normalized class
@@
This class satisfies a reciprocity law which directly relates it to the $p$-adic $L$-function $L_{p}({\mathbf{f}})$, thereby avoiding the inconvenient appearance of the adjoint $p$-adic $L$-function and the auxiliary Kitagawa–Mazur $p$-adic $L$-function. However, as we later discuss, we must exercise more caution in the Eisenstein setting, since both $p$-adic $L$-functions vanish at the critical point.
@@ 5. Deformation of Beilinson–Kato elements
The aim of this section is studying the two-variable Beilinson–Kato elements in the case where they pass through a critical Eisenstein family. In particular, we want to relate the natural projection of that cohomology class with the circular units discussed in Section 3. Since the aim of this section is establishing Theorem 1.3, we assume that $r$ is even, although many of the results also hold in the odd case.
5.1. Set-up for the generation process
@@ 5.2. Families over punctured discs
with the following interpolation property:
  • If $k\geq 0$ is an integer with $k\neq r$, then we have
@@
The result follows verbatim the reasoning of [LZ16, Thm. 5.4.2], which establishes an analogous result when all integer-weight specializations of ${\mathbf{f}}$ are classical. In this case, the result does not apply for $k=r$, but inverting $X$ resolves the issue. ∎
@@ 5.3. The $p$-adic $L$-function and a factorization formula
An equivalent definition for $\log_{p}^{[k]}$ is $\log_{p}^{[k]}(\sigma)=w(\sigma)(w(\sigma)-1)\cdots(w(\sigma)-k+1)$, where $w(\sigma)=\frac{\log_{p}(\sigma(1+p))}{\log_{p}(1+p)}$. The analytic function $\log_{p}^{[k]}$ vanishes to order 1 at the characters $\sigma$ of the form $z\mapsto z^{j}\chi(z)$, where $\chi$ is a finite order Dirichlet character and $j$ is an integer such that $10\leq j\leq k-$ 1.
@@
Following the analogy with the Beilinson–Flach case, we would expect the class $\kappa(f_{\beta})$ to be divisible by the logarithmic distribution $\log^{[r+1]}\in\mathcal{H}_{\Gamma}$. Let us sketch why the same argument that we had used in that case does not work here. The specialization of $\kappa({\mathbf{f}})$ at a locally-algebraic character of $\Gamma$ of degree $j\in\{0,\dots,r\}$ factors through the image of $\mathscr{D}_{U-j}\otimes\operatorname{TSym}^{j}$ in $\mathscr{D}_{U-(r+1)}\otimes\operatorname{TSym}^{(r+1)}$, and the maps $\Pr_{{\mathbf{f}}}^{[j]}$ and $\Pr_{{\mathbf{f}}}^{[r+1]}$ agree on this image up to a non-zero scalar. Since the $\Pr_{{\mathbf{f}}}^{[j]}$ for $0\leqslant j\leqslant r$ do not have poles at $X=0$, it follows that the specializations of $\kappa({\mathbf{f}})$ at a pair $(r,\chi)$, for $\ell\geqslant r$ and $\chi$ locally-algebraic of degree $\in\{0,\dots,r\}$, interpolate the projections of the classical Beilinson–Kato classes to the $(E_{r+2}^{\operatorname{crit}})$-eigenspaces in classical cohomology. However, in this setting we cannot assure whether or not these classes arise as suitable projections of classes in the cohomology of $X_{1}(N)$, that we know that there are always 0. However, we will get a similar conclusion regarding the divisibility by a logarithmic term using the $p$-adic $L$-function.
Consider now the two-variable $p$-adic $L$-function $L_{p}({\mathbf{f}})$ defined over $U\times\mathcal{W}$, as introduced in Section 4.1. For this construction, the interpolation property applies also at $k=r$ without any further modifications; and here the complex $L$-function factors, via Artin formalism, as
@@
See [BD15, §5, 6]. ∎
More concretely, the proof of the first statement relies on the application $\Theta_{k}$, defined in [Bel11], which allows us to relate the critical weight $k$ Eisenstein series with the ordinary Eisenstein series of weight $-2-k$. Furthermore, the factor $L_{p}(\tau,s-k+1)$ is the image under the Coleman map of the system of circular units, thus providing a natural link with the other Euler system explored in this note.
@@
This follows by taking derivatives along the weight direction in Proposition 4.4. Then, specializing at $k=r$ and applying Theorem 5.6, the result follows. ∎
@@ 5.4. Perrin-Riou maps
$\operatorname{Col}_{\mathbf{a}_{{\mathbf{f}}}^{-}}=\left\langle\mathcal{L}^{\operatorname{PR}}_{\mathcal{F}^{-}V({\mathbf{f}})^{*}}(-),\mathbf{a}_{{\mathbf{f}}}^{-}\right\rangle:H^{1}_{\mathrm{Iw}}(\mathbf{Q}_{p,\infty},\mathcal{F}^{-}D({\mathbf{f}})^{*})\to\mathcal{O}(U\times\mathcal{W})$
@@
where $e_{n}$ is the standard basis of $\mathbf{Z}_{p}(n)$ and $d_{r}$ is a non-zero constant. Since multiplication by $t^{r+1}$ corresponds to multiplication by $\log^{[r+1]}$ on the $\mathcal{H}_{\Gamma}$ side (see e.g. [LLZ11]), we conclude that
$\operatorname{Col}_{\mathbf{a}_{{\mathbf{f}}}^{-}}(\kappa({\mathbf{f}}))\bmod X=d_{r}\left\langle\mathcal{L}^{\text{PR}}_{\tau^{-1}}(\kappa(f_{\beta})),\omega_{f_{k}}\right\rangle.$
For stating the following result, we recover the setting of Section 4.3. In particular, we are working with two-variables $(k,s)$, which may be understood in terms of the variables $(k,\ell)$ corresponding to the weight specializations of the Coleman family ${\mathbf{f}}$ and the ordinary Eisenstein family ${\mathbf{E}}$, where, with the notations of Section 4.3, $(\chi_{1},\chi_{2})=(1,\chi)$. In order to simplify the arguments in the next result, we assume that $L_{p}^{\prime}(f_{\beta},\chi)\neq 0$, although a minor modification of the leading term argument solves the issue.
@@
Under the current assumptions, if $L_{p}^{\prime}(f_{\beta},\chi)\neq 0$, we have
@@
where $d_{{\mathbf{f}}}({\mathbf{k}})$ is a meromorphic function on $U$ alone, regular and non-vanishing at all integer weights $k\geqslant-1$ except possibly at $r$ itself, and regular at $k=r$. Recall that $\mathcal{E}_{N}$ is the product of the bad Euler factors introduced in Definition 4.7.
@@
It follows easily from the reciprocity laws for Rankin–Selberg convolutions (of cuspidal forms) that the quotient $\operatorname{Col}_{\mathbf{a}_{{\mathbf{f}}}^{-}}(\kappa({\mathbf{f}}))/(\mathcal{E}_{N}\cdot L_{p}({\mathbf{f}},{\mathbf{E}}))$ is a function of ${\mathbf{k}}$ alone, and this ratio does not vanish at any integer $k\geqslant-1$ where $f_{k}$ is classical; it is equal to the fudge-factor $d_{k}$ defined above.
Moreover, since $L_{p}({\mathbf{f}},{\mathbf{E}})$ is well-defined and non-zero along $\{r\}\times\mathcal{W}$, we conclude that $d_{{\mathbf{f}}}({\mathbf{k}})$ does not have a pole at ${\mathbf{k}}$ (although it might have a zero there). ∎
@@
The non-vanishing of $L_{p}^{\prime}(f_{\beta},\chi)$ must be understood as a very mild assumption, since the character $\chi$ can be chosen among all the possible Dirichlet characters satisfying a parity constraint. Further, if $L_{p}^{\prime}(f_{\beta},\chi)=0$, the function $d_{{\mathbf{f}}}({\mathbf{k}})$ has a (not necessarily simple) pole; in that case, our next results can be adapted considering the leading term of $d_{{\mathbf{f}}}({\mathbf{k}})$ in the Laurent expansion.

### v3

@@ 4. Beilinson–Kato elements
The Kato Euler system (or Beilinson–Kato Euler system) is one of the most significant constructions in the study of $p$-adic methods in Number Theory. It has enabled the proof of new cases of the Bloch–Kato conjecture and provided a general approach to establish one divisibility in the Iwasawa main conjecture. More concretely, the Kato Euler system of [Kat04] can be attached to a general Coleman family $\mathbf{f}$}, as discussed e.g.  by Hansen in [Han15] or by Benois and Büyükboduk in [BB22] and [BB24]; for the the case of Hida families, the theory is known since the work of Ochiai [Och03]. In this section, we begin by recalling the (two-variable) $p$-adic $L$-function of Kitagawa and Mazur, and later discuss the reciprocity law, emphasizing the role of the (one-variable) adjoint $p$-adic $L$-function. This discussion, and in particular the proof of the reciprocity law, does not hold at the critical Eisenstein point, so we later need to carry out certain modifications to adapt it.
@@ 4.1. The Kitagawa–Mazur $p$-adic $L$-function
Let $f_{k}$ be the newform associated to the weight $k+2$ specialization of $\mathbf{f}$}, for some $k\geq 0$. Assume that $f_{k}$ does not correspond to a $p$-stabilization of an Eisenstein series, and let $\mathbf{Q}(f_{k})$ the field of coefficients of $f_{k}$. Then, for each sign $\pm$, the eigenspace in Betti cohomology of $Y_{1}(N)$ on which the Hecke operators act by the eigensystem of $f_{k}$ is one-dimensional, and we choose bases $\gamma_{f}^{\pm}$ of theses spaces. These determine complex periods $\Omega_{f_{k}}^{\pm}\in\mathbf{C}^{\times}$.
With the previous notations, let $U=\mathrm{Spa}(A_{U},A_{U}^{\circ})\subset\mathcal{C}^{0}$ be the affinoid open over which the family $\mathbf{f}$} is defined. Then we have an $A_{U}$-module $V({\mathbf{f}})^{*}$, free of rank 2. We choose bases $\gamma^{\pm}_{{\mathbf{f}}}$ of the complex-conjugation eigenspaces $V({\mathbf{f}})^{(c=\pm 1)}$, which are free of rank one over $A_{U}$. As pointed out in [LRS24, §B2.1], we need to extend the definition of $\Omega_{f_{k}}^{\pm}$ in such a way that they lie in the space $(L\otimes_{\mathbf{Q}(f_{k})}\mathbf{C})^{\times}$.
@@
$L_{p}({\mathbf{f}})\in A_{U}\hat{\otimes}_{\mathbf{Q}_{p}}\mathcal{H}_{\Gamma}$
@@
More precisely, the value at $(k,j)$, with $0\leqslant j\leqslant k-2$, interpolates the $L$-value $L(f_{k},1+j)$ up to appropriate factors, where $f$ is the weight $k$ specialization. See e.g. [FK24, §4] or [BB22] for the precise form of the interpolation property, which involves different kinds of factors.
@@ 4.2. The adjoint $p$-adic $L$-function
For this section, we mainly follow the works of Hida [Hid16] and Bellaïche [Bel21, §9], who constructed an adjoint $p$-adic $L$-function, using Kim’s scalar product on overconvergent modular symbols. This was later generalized to a more general setting in the recent works of Balasubramanyam, Bergdall, and Longo [BBL20], or Lee and Wu [LW25]. See also [Mak23] for a detailed treatment of the topic.
@@
As pointed out in [Mak23], Bellaïche’s interpolation formula is only given at crystalline points of trivial nebentype, while in Thm. 2.3.7 of loc. cit. the author gives a much more general result which, however, is only valid in the setting of ordinary Hida families.
@@ 4.3. Hida–Rankin vs Kitagawa–Mazur
This follows by a direct comparison of the critical values, as in [BD14, Thm. 3.4], noting that the Euler factors also match. Observe that in loc. cit. the authors restrict to the ordinary case, while e.g. in [BDV22, §4.3] the proof is adapted to general families.
Observe, however, that in this case we cannot discard the factors contributing to the adjoint, since they can be zero (and indeed, they vanish in the situations we want to study). ∎
@@
Remark 4.5.
We do not actually need the existence of the adjoint $p$-adic $L$-function in this setting. We can just claim an equality of the form
$L_{p}({\mathbf{f}},{\mathbf{g}})(k,\ell,s)\times\mathfrak{g}(k)=L_{p}({\mathbf{f}},\chi_{1})(k,s)\times L_{p}({\mathbf{f}},\chi_{2})(k,s-\ell+1),$
where $\mathfrak{g}(k)$ is a $p$-adic function interpolating the adjoint periods.
@@ 4.4. Kato’s Euler system
Recall that we have chosen bases $\gamma^{\pm}_{{\mathbf{f}}}$ of the complex-conjugation eigenspaces $V({\mathbf{f}})^{(c=\pm 1)}$, which are free of rank one over $A_{U}$. For the moment, we assume that we are outside the Eisenstein case.
@@
$\mathcal{E}_{N}(x)=\prod_{\ell\mid N}(1-a_{\ell}({\mathbf{f}})\sigma_{\ell}^{-1})\in A_{U}\hat{\otimes}_{\mathbf{Q}_{p}}\mathcal{H}_{\Gamma}.$
@@
More precisely, let $\mathcal{F}^{+}V({\mathbf{f}})$ denote the rank 1 unramified subrepresentation of $V({\mathbf{f}})$ over $A_{U}$ (which exists since $\mathbf{f}$} is ordinary); and let $\eta_{{\mathbf{f}}}\in\mathbf{D}_{\mathrm{cris}}(\mathcal{F}^{+}V({\mathbf{f}}))$ be the canonical vector constructed in [KLZ17], which is characterized by interpolating the classes $\eta_{f}$ of Section 2.3 for each classical specialization $f$ of $\mathbf{f}$}. Then we have the following reciprocity law, where we have an extra contribution coming from the Euler factors at bad primes. For the following proposition, write $\mathcal{L}_{\mathcal{F}^{-}V({\mathbf{f}})^{*}}^{\operatorname{PR}}$ for the Perrin-Riou logarithm attached to $\mathcal{F}^{-}V({\mathbf{f}})^{*}$.
We state the reciprocity law in the framework of [BB22], where we avoid the critical Eisenstein points.
@@
Let $\mathbf{f}$} be a Coleman family such that all its specializations are cuspidal (and which, in particular, does not pass through the critical $p$-stabilization of an Eisenstein series). Then, the following explicit reciprocity law holds:
@@
This is Kato’s explicit reciprocity law in the framework of Coleman families, as developed e.g. in [BB22]. Observe that the presence of the adjoint $p$-adic $L$-function, and also the auxiliary Kitagawa–Mazur $p$-adic $L$-function $L_{p}({\mathbf{f}},\chi)$ is usually treated as a unique contribution in terms of a non-vanishing factor, but we want to emphasize this dependence here since we will later need to be more circumspect around this point. ∎
Remark 4.10.
There are some crucial differences between the theory of Beilinson–Flach and Beilinson–Kato classes. In particular, the latter depends on the choice of a suitable class in homology; the results of Ash and Stevens guarantee that this can be done in such a way that the Kato class is non-zero. Observe, however, that in the arguments of [BB22] it is crucial that the critical point is cuspidal.
Proposition 4.9 must be understood as an equality in two variables, with $L_{p}({\mathbf{f}},\chi)$ playing an auxiliary role. In many case, it is possible to invert the factor $L_{p}({\mathbf{f}},\chi)$ and normalize the Kato class dividing by it; however, in the Eisenstein case one must be more circumspect and we cannot do this normalization.
@@
In the setting where $\mathbf{f}$} is not a critical Coleman family, the classes are usually renormalized to avoid dependence on the character $\chi$. More precisely, we may consider the normalized class
@@
This class satisfies a reciprocity law that directly relates it to the $p$ -adic $L$ -function $L_{p}({\mathbf{f}})$, thus avoiding the inconvenient appearance of the adjoint $p$-adic $L$-function and the auxiliary Kitagawa–Mazur $p$-adic $L$-function. However, as we later discuss, we must exercise more caution in the Eisenstein setting, since both $p$-adic $L$-functions vanish at the critical point.
@@ 5. Deformation of Beilinson–Kato elements
The aim of this section is to study the two–variable Beilinson–Kato elements in the case where they pass through a critical Eisenstein family. In particular, we want to relate the natural projection of that cohomology class with the circular units discussed in Section 3. Since the aim of this section is to establish Theorem 1.3, we assume that $r$ is even, although many of the results also hold in the odd case.
In particular, as we have already emphasized, the results of Benois and Büyükboduk do not hold in our case, and we need to adapt both the interpolation of the Beilinson–Kato class and the explicit reciprocity law.
5.1. Set-up for the degeneration process
@@ 5.2. Families over punctured discs
with the following interpolation property: if $k\geq 0$ is an integer with $k\neq r$, then we have
@@
To begin with, observe that the results of [BB22] do not apply in this setting, since, in particular, the key Prop. 6.5 of loc. cit. requires all the specializations being cuspidal, as it occurs in the Beilinson–Flach case in [LZ16].
The result follows verbatim the reasoning of [LZ16, Thm. 5.4.2], which establishes an analogous result when all integer-weight specializations of $\mathbf{f}$} are classical. In this case, the result does not apply for $k=r$, but inverting $X$ resolves the issue. In particular, the analysis of the comparison maps in cohomology developed in [LRS24, §A] makes clear that we are only introducing a simple pole at the critical Eisenstein point. ∎
@@ 5.3. The $p$-adic $L$-function and a factorization formula
An equivalent definition for $\log_{p}^{[k]}$ is $\log_{p}^{[k]}(\sigma)=w(\sigma)(w(\sigma)-1)\cdots(w(\sigma)-k+1)$, where $w(\sigma)=\frac{\log_{p}(\sigma(1+p))}{\log_{p}(1+p)}$. The analytic function $\log_{p}^{[k]}$ vanishes to order 1 at the characters $\sigma$ of the form $z\mapsto z^{j}\chi(z)$, where $\chi$ is a finite order Dirichlet character and $j$ is an integer such that $0\leq j\leq k-1$.
@@
Following the analogy with the Beilinson–Flach case, we would expect the class $\kappa(f_{\beta})$ to be divisible by the logarithmic distribution $\log^{[r+1]}\in\mathcal{H}_{\Gamma}$. Let us sketch why the same argument that we had used in that case does not work here. The specialization of $\kappa({\mathbf{f}})$ at a locally-algebraic character of $\Gamma$ of degree $j\in\{0,\dots,r\}$ factors through the image of $\mathscr{D}_{\lambda_{V}-j}^{s}\otimes\operatorname{TSym}^{j}$ in $\mathscr{D}_{\lambda_{V}-(r+1)}^{s}\otimes\operatorname{TSym}^{(r+1)}$, and the maps $\Pr_{{\mathbf{f}}}^{[j]}$ and $\Pr_{{\mathbf{f}}}^{[r+1]}$ agree on this image up to a non-zero scalar. Since the $\Pr_{{\mathbf{f}}}^{[j]}$ for $0\leqslant j\leqslant r$ do not have poles at $X=0$, it follows that the specializations of $\kappa({\mathbf{f}})$ at a pair $(r,\chi)$, for $\ell\geqslant r$ and $\chi$ locally-algebraic of degree $\in\{0,\dots,r\}$, interpolate the projections of the classical Beilinson–Kato classes to the $(E_{r+2}^{\operatorname{crit}})$-eigenspaces in classical cohomology. However, in this setting, we cannot ensure whether or not these classes arise as suitable projections of classes in the cohomology of $X_{1}(N)$.
Consider now the two-variable $p$-adic $L$-function $L_{p}({\mathbf{f}})\in A_{V}\hat{\otimes}_{\mathbf{Q}_{p}}\mathcal{H}_{\Gamma}$ as introduced in Section 4.1. For this construction, the interpolation property applies also at $k=r$ without any further modifications; and here the complex $L$-function factors, via Artin formalism, as
@@
See [BD15, §5, 6]. More concretely, the proof of the first statement relies on the application $\Theta_{k}$, defined in [Bel11], which allows us to relate the critical weight $k$ Eisenstein series with the ordinary Eisenstein series of weight 2k-2-k. ∎
Furthermore, the factor $L_{p}(\tau,s-k+1)$ is the image under the Coleman map of the system of circular units, thus providing a natural link with the other Euler system explored in this note.
@@
This follows by taking derivatives along the weight direction in Proposition 4.4, and using that both $L_{p}^{\prime}(\operatorname{Ad}{\mathbf{f}})$ and $L_{p}({\mathbf{f}},\chi)$ vanish at the critical Eisenstein point. Then, specializing at $k=r$ and applying Theorem 5.6, the result follows. ∎
@@ 5.4. Perrin-Riou maps
At this point, and for establishing an explicit reciprocity law, the theory presents notable differences with that of Beilinson–Flach elements. In particular, the Beilinson–Kato element depends on the choice of a homology period. By a result of Ash–Stevens [Kat04, Thm. 13.6] in the cuspidal case, there exists a choice of periods which guarantees that the class is non-zero. Further, the cohomology class obtained for a different choice of periods gives another class that, for any specialization, is a multiple of the one coming from Ash–Stevens. Unfortunately, in the Eisenstein case, we cannot claim that such a good normalization exists, so it is possible that all the possible choices give the zero class. Then, our results of this section must be interpreted as follows: if there exists a choice of period for which $\kappa({\mathbf{f}})$ is not zero, we prove a reciprocity law for that class, which later allows us to relate it with a circular unit; if not, we just get the zero class.
Remark 5.8.
This discussion suggests, as a very natural question, if we can prove that there exists a normalization of the Kato class such that $\kappa({\mathbf{f}})$ is non-zero.
@@
$\operatorname{Col}_{\mathbf{a}_{{\mathbf{f}}}^{-}}=\left\langle\mathcal{L}^{\operatorname{PR}}_{\mathcal{F}^{-}V({\mathbf{f}})^{*}}(-),\mathbf{a}_{{\mathbf{f}}}^{-}\right\rangle:H^{1}_{\mathrm{Iw}}(\mathbf{Q}_{p,\infty},\mathcal{F}^{-}D({\mathbf{f}})^{*})\to\mathcal{H}_{\Gamma}\hat{\otimes}_{\mathbf{Q}_{p}}A_{U}$
@@
where $e_{n}$ is the standard basis of $\mathbf{Z}_{p}(n)$ and $d_{r}$ is a constant. Since multiplication by $t^{r+1}$ corresponds to multiplication by $\log^{[r+1]}$ on the $\mathcal{H}_{\Gamma}$ side (see e.g. [LLZ11]), we conclude that
$\operatorname{Col}_{\mathbf{a}_{{\mathbf{f}}}^{-}}(\kappa({\mathbf{f}}))\bmod X=d_{r}\cdot\log^{[r+1]}\cdot\left\langle\mathcal{L}^{\text{PR}}_{\tau^{-1}}(\kappa(f_{\beta})),\omega_{f_{k}}\right\rangle.$
We now establish the explicit reciprocity law for a family $\mathbf{f}$} through a critical Eisenstein point. Note that this is not automatic from our previous result, which allows us to claim it in a neighborhood of the critical Eisenstein point, thus leading to the present of some annoying constants.
In order to state the next result, we recover the setting of Section 4.3. In particular, we are working with two-variables $(k,s)$, which may be understood in terms of the variables $(k,\ell)$ corresponding to the weight specializations of the Coleman family $\mathbf{f}$} and the ordinary Eisenstein family $\mathbf{E}$}, where, with the notations of Section 4.3, $(\chi_{1},\chi_{2})=(1,\chi)$. In order to simplify the arguments in the next result, we assume that $L_{p}^{\prime}(f_{\beta},\chi)\neq 0$, although a minor modification of the leading term argument can solve the issue.
@@
Under Assumption 1.1 and Assumption 1.2, we have
@@
where $d_{{\mathbf{f}}}({\mathbf{k}})$ is either the zero function, or, if not, it is a meromorphic function on $U$ alone, regular and non-vanishing at all integer weights $k\geqslant-1$ except possibly at $r$ itself, and regular at $k=r$. Recall that $\mathcal{E}_{N}$ is the product of the bad Euler factors introduced in Definition 4.8.
@@
It follows from the reciprocity laws for Rankin–Selberg convolutions (of cuspidal forms) that the quotient $\operatorname{Col}_{\mathbf{a}_{{\mathbf{f}}}^{-}}(\kappa({\mathbf{f}}))/(\mathcal{E}_{N}\cdot L_{p}({\mathbf{f}},{\mathbf{E}}))$ is a function of $\mathbf{k}$} alone, and this ratio is either identically zero or, following the discussion of [BB22, §7], it does not vanish at any integer $k\geqslant-1$ where $f_{k}$ is classical; it is equal to the fudge-factor $d_{k}$ defined above.
Moreover, since under the current assumptions $L_{p}({\mathbf{f}},{\mathbf{E}})$ is well defined and non-zero along $\{r\}\times\mathcal{W}$, we conclude that $d_{{\mathbf{f}}}({\mathbf{k}})$ does not have a pole at $\mathbf{k}$} (although it might have a zero there). ∎
@@
The non-vanishing of $L_{p}^{\prime}(f_{\beta},\chi)$ must be understood as a very mild assumption, since the character $\chi$ can be chosen among all the possible Dirichlet characters satisfying a parity constraint. Further, if $L_{p}^{\prime}(f_{\beta},\chi)=0$, the function $d_{{\mathbf{f}}}({\mathbf{k}})$ can have a (not necessarily simple) pole; in that case, our results can be adapted considering the leading term of $d_{{\mathbf{f}}}({\mathbf{k}})$ in its Laurent expansion.